# Ratio asymptotics and zero density for orthogonal polynomials with varying Verblunsky coefficients

### v1

@@ Abstract.
We study asymptotic behavior of orthogonal polynomials on the unit circle with varying Verblunsky coefficients $\alpha_{n,N}$.
First, we give a streamlined proof of ratio asymptotics for orthogonal and paraorthogonal polynomials in the case when $\alpha_{n,N}$ is asymptotically constant or asymptotically periodic as $n/N\to t>0$.
Second, we determine the asymptotic zero density of paraorthogonal polynomials in the locally constant and locally periodic regimes, and obtain analogous results for orthogonal polynomials under an additional condition $|\alpha_{n-1,N}|^{1/n}\to 1$.
@@ 1. Introduction
In this paper we investigate orthogonal polynomials $\Phi_{n,N}$ on the unit circle determined by Verblunsky coefficients $\alpha_{n,N}$ that are varying with an additional index $N\in\mathbb{N}$. In other words, for each $N\in\mathbb{N}$, the polynomial $\Phi_{n,N}$ is monic of degree $n$ and satisfies the Szegő recurrence
$\Phi_{n+1}(z)=z\Phi_{n}(z)-\bar{\alpha}_{n}\Phi_{n}^{*}(z),\quad n\geq 0,$
with the initial condition $\Phi_{0,N}(z)=1$ for all $N\in\mathbb{N}$. Here and throughout, the bar denotes complex conjugation and $\Phi_{n,N}^{*}(z)=z^{n}\overline{\Phi_{n,N}\left(1/\bar{z}\right)}$ is the reversed polynomial.
A closely related concept is that of paraorthogonal polynomials, which play an important role in approximation theory and are defined by
$\Phi^{(\beta)}_{n+1,N}(z)=z\Phi_{n,N}(z)-\bar{\beta}\Phi_{n,N}^{*}(z),\quad n\geq 0,$ (1.1)
where the parameter $\beta$ belongs to the unit circle ${\partial\mathbb{D}}\equiv\{z\in{\mathbb{C}}\mid|z|=1\}$.
We are interested in the behavior of $\Phi_{n,N}$, $\Phi^{*}_{n,N}$, $\Phi_{n,N}^{(\beta)}$, and their zeros when both $n$ and $N$ tend to infinity, but in such a way that the ratio $n/N\to t$, where $t>0$ is a fixed constant.
In our first result, Theorem 5, we establish the ratio asymptotics of $\Phi_{n,N}$, $\Phi^{*}_{n,N}$, $\Phi_{n,N}^{(\beta)}$, as $n/N\to t$, assuming the varying Nevai–López condition
$\lim_{n/N\to s}\alpha_{n,N}=\alpha.$ (1.2)
While this result is not surprising – analogous statements are well known in the non-varying case $\alpha_{n,N}\equiv\alpha_{n}$ (see  [BarLop, Khr02] and [simon_opuc1, simon_opuc2]) – our approach via the operator theory of CMV matrices offers a streamlined proof that may be of interest even in the classical setting. In particular, it allows for an explicit identification of the limiting ratio in terms of classical objects such as the Schur, Carathéodory, and $m$-functions, and it establishes convergence on arbitrary compact subsets of $\mathbb{C}$}, provided one avoids the limiting zero set of the polynomials.
Furthermore, the exact same approach extends without any difficulty to much wider generality, for example to asymptotically periodic setting
@@
for some period $p\in\mathbb{N}$. This constitutes our second result, Theorem 7. In the non-varying case, it improves upon the result of Barrios and López [BarLop, Thm. 2], which determined the limit $\Phi_{n+p}(z)/\Phi_{n}(z)$, cf. Corollary 9.
Our next result concerns the asymptotics of the zero-counting measures
$\nu_{n,N}$ $=\frac{1}{n}\sum_{j=1}^{n}\delta_{z_{j,n,N}},$ (1.4)
$\nu^{(\beta)}_{n,N}$ $=\frac{1}{n}\sum_{j=1}^{n}\delta_{z^{(\beta)}_{j,n,N}}$, (1.5)
where $z_{j,n,N}$ and $z^{(\beta)}_{j,n,N}$ are the zeros of $\Phi_{n,N}$ and $\Phi_{n,N}^{(\beta)}$, respectively, counted with multiplicity.
It is clear that the mere conditions (1.2) or (1.3) are generally insufficient to guarantee the existence of a weak limit for these measures, as they constrain only a small portion of the Verblunsky coefficients. This leads us to the so-called asymptotically locally constant setting, in which
@@
A model situation for this setting is obtained by sampling a piece-wise continuous function $\alpha(s)$ on a regular grid: $\alpha_{n,N}=\alpha\left(n/N\right)$ for $n\leq tN$ (and, for instance, $\alpha_{n,N}=0$ for $n>tN$).
@@
Our Theorem 10 states that under (1.6), the weak limit of the zero-counting measure $\nu^{(\beta)}_{n,N}$ of paraorthogonal polynomials exists and is equal to
@@
where the measure $\nu_{|\alpha(s)|}$ is the equilibrium measure of the arc $\Gamma_{|\alpha(s)|}$. If we additionally assume that $|\alpha_{n-1,N}|^{1/n}\to 1$ then Theorem 18 yields the same weak limit for the zero-counting measure $\nu_{n,N}$ of orthogonal polynomials.
@@
Finally, the analogous result holds (see Theorem 14 and Remark 20) in the asymptotically locally periodic setting, which is
$\lim_{kp/N\to s}\alpha_{kp+j,N}=\alpha_{j}(s),\,\mbox{ for {a.e.} }s\in[0,t],$ (1.8)
@@
where $\alpha_{1}(s),\ldots,\alpha_{p}(s)$ are $p$ integrable functions on $[0,t]$. In this case, the measure $\nu_{|\alpha(s)|}$ appearing on the right-hand side of (1.7) is replaced by the equilibrium measure of a certain finite-gap set.
It is worth mentioning that the condition (1.6) appears to be very natural from the perspective of discretizing Krein systems and one-dimensional Dirac operators, see [denisov_imrs06, Kre55]. The problem of understanding asymptotic zero densities in the varying setting also is also closely related to questions in random matrix theory concerning global eigenvalue distribution of orthogonal polynomial ensembles on the unit circle, see [BreOfn25, DuiKoz].
There is a vast literature on the ratio asymptotics and zero asymptotics of orthogonal polynomials on the unit circle. The most relevant results that we use are in [BarLop, golinski_maa99, Khr02, mha-saf_jat90, Sim09], as well as [simon_opuc2, Sec. 8].
@@
For orthogonal polynomials on the real line, the asymptotic zero distribution with varying Jacobi parameters was analyzed by Kuijlaars and Van Assche [kuijlaarsvanassche_jat99]. Related results had already appeared in the Szegő-type limit formulas for Hermitian variable Toeplitz matrices due to Kac, Murdock and Szegő [kms_jrma53], later revisited by Tilli [tilli_laa98]; see also [pastur_mps96, bourget-etal_jjm18]. These works also influenced the development of locally Toeplitz sequences and their applications in numerical analysis [garoni-serracapizzano_vol1, garoni-serracapizzano_vol2].
@@
On the other hand, the article [kuijlaarsvanassche_jat99] had a significant impact on the orthogonal polynomial community. Following its appearance, several generalizations and extensions were developed: to the asymptotically periodic case in [vanassche_ss99], to discontinuous limiting functions in [kuijlaars-serracapizzano_jat01], to Laurent orthogonal polynomials with varying recurrence coefficients in [coussement-vanassche_jpa05], and to variable multiple orthogonal polynomials in [coussement-etal_tams08]. Our results cover the full range of [kuijlaarsvanassche_jat99, vanassche_ss99, kuijlaars-serracapizzano_jat01] in the unit-circle setting, including both asymptotically periodic behavior and discontinuous sampling functions.
Throughout the paper, we use the abbreviations OPUC and POPUC for orthogonal polynomials on the unit circle and paraorthogonal polynomials on the unit circle, respectively.
@@
The paper is organized as follows. Section 2 collects notation and the necessary background from the theory of orthogonal polynomials and potential theory.
@@
Section 3 derives ratio asymptotics for varying POPUC and OPUC: the asymptotically constant case is treated in Section 3.1, and the asymptotically periodic case in Section 3.2.
@@
Section 4 establishes the asymptotic zero distribution for POPUC with varying Verblunsky coefficients. We present two proofs of the main result: one combining ratio asymptotics with a potential-theoretic argument (Section 4.1), and another based on the method of moments (Section 4.2). The asymptotically periodic case is discussed in Section 4.3.
@@
Section 5 concerns the asymptotic zero distribution for OPUC with varying Verblunsky coefficients. The connection between POPUC and OPUC is handled via balayage in Section 5.1, and the main result (under an additional assumption) appears in Section 5.2.
@@
Finally, Section 6 presents several illustrative examples where the limiting densities can be computed explicitly or expressed in terms of classical special functions, together with numerical plots of these densities.
@@ 2. Preliminaries
Throughout the paper we use the following notation: $\overline{{\mathbb{C}}}={\mathbb{C}}\cup\{\infty\}$ for the Riemann sphere; $\mathbb{N}=\{1,2,\dots\}$ for positive integers; $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ for non-negative integers; ${\mathbb{D}}=\{z\in{\mathbb{C}}\mid|z|<1\}$ for the open unit disk; $\overline{{\mathbb{D}}}=\{z\in{\mathbb{C}}\mid|z|\leq 1\}$ for the closed unit disk; ${\partial\mathbb{D}}=\{z\in{\mathbb{C}}\mid|z|=1\}$ for the unit circle; ${\mathbb{D}}^{\infty}=\{\{\alpha_{n}\}_{n\in\mathbb{N}_{0}}\mid\alpha_{n}\in{\mathbb{D}}\mbox{ for all }n\in\mathbb{N}_{0}\}$ for the set of sequences with elements from $\mathbb{D}$}; $\ell^{2}(\mathbb{N}_{0})=\{\{x_{n}\}_{n=0}^{\infty}\mid\sum_{n=0}^{\infty}|x_{n}|^{2}<\infty\}$ for the Hilbert space of complex square summable sequences; $\langle\cdot,\cdot\rangle$ for the inner product on $\ell^{2}(\mathbb{N}_{0})$ linear in the second argument; and $e_{i}$ for the $i$-th vector of the standard basis of either ${\mathbb{C}}^{n}$ or $\ell^{2}(\mathbb{N}_{0})$ which is always clear from the context, i.e., $(e_{i})_{i}=1$ and $(e_{i})_{j}=0$ for $j\neq i$.
@@
$\{n_{j}\}_{j\in\mathbb{N}},\{N_{j}\}_{j\in\mathbb{N}}\subseteq\mathbb{N},$ (2.1)
$n_{j},N_{j}\to\infty\mbox{ and }n_{j}/N_{j}\to t,\mbox{ as }j\to\infty.$
@@
expresses the limit of the doubly indexed sequence of measures $\mu_{n,N}$ converging to $\mu$ weakly, i.e., in the weak topology. Similarly,
$\operatorname*{s-lim\ }_{n/N\to t}A_{n,N}=A$
stands for the limit of the doubly indexed sequence of bounded operators $A_{n,N}$ converging to $A$ strongly.
@@ 2.1. Orthogonal and paraorthogonal polynomials
Given a non-trivial probability measure $\mu$ on $\partial\mathbb{D}$} (i.e, one with infinite support), let $\Phi_{n}$ denote the monic orthogonal polynomial on the unit circle (OPUC), which is the unique monic polynomial of degree $n$ satisfying the orthogonality relation
$\int_{0}^{2\pi}\Phi_{n}(e^{{\rm i}\theta})e^{-{\rm i}k\theta}\,{\rm d}\mu(e^{{\rm i}\theta})=0,\quad\mbox{for all }k=0,\ldots,n-1.$ (2.2)
These polynomials satisfy the Szegő recursion relations
@@
If the Verblunsky coefficient $\alpha_{n-1}$ is replaced by a unimodular parameter $\beta\in{\partial\mathbb{D}}$ one obtains the so-called paraorthogonal polynomials on the unit circle (POPUC)

### v2

@@ Abstract.
We study asymptotic behavior of orthogonal polynomials on the unit circle with varying Verblunsky coefficients $\alpha_{n,N}$ when the ratio $n/N$ converges as $n,N\to\infty$. First, we give a streamlined proof of ratio asymptotics for orthogonal and paraorthogonal polynomials in the case of asymptotically constant and asymptotically periodic coefficients $\alpha_{n,N}$. Second, we determine the asymptotic zero distribution of paraorthogonal polynomials in the locally constant and locally periodic regimes. Analogous results are obtained for orthogonal polynomials under a mild additional condition on the varying coefficients.
@@ 1. Introduction
We investigate orthogonal polynomials $\Phi_{n,N}$ on the unit circle determined by Verblunsky coefficients $\alpha_{n,N}$ that are varying with an additional index $N\in\mathbb{N}$. In other words, given a doubly-indexed sequence $\alpha_{n,N}$ from the unit disk ${\mathbb{D}}\equiv\{z\in{\mathbb{C}}\mid|z|<1\}$, the monic polynomial $\Phi_{n,N}$ of degree $n$ is generated by the Szegő recurrence
$\Phi_{n+1,N}(z)=z\Phi_{n,N}(z)-\bar{\alpha}_{n,N}\Phi_{n,N}^{*}(z),\quad n\geq 0,$
with the initial condition $\Phi_{0,N}(z)=1$ for all $N\in\mathbb{N}$. Here and throughout, the bar denotes complex conjugation and $\Phi_{n,N}^{*}(z)=z^{n}\,\overline{\Phi_{n,N}\left(1/\bar{z}\right)}$ is the reversed polynomial. Given an additional parameter $\beta$ in the unit circle ${\partial\mathbb{D}}\equiv\{z\in{\mathbb{C}}\mid|z|=1\}$, the closely related concept of paraorthogonal polynomials $\Phi_{n,N}^{(\beta)}$ is, for each $N\in\mathbb{N}$, defined by
$\Phi^{(\beta)}_{n+1,N}(z)=z\Phi_{n,N}(z)-\bar{\beta}\Phi_{n,N}^{*}(z),\quad n\geq 0.$
In this article, we study the asymptotic behavior of the polynomials $\Phi_{n,N}$, $\Phi_{n,N}^{(\beta)}$, and of their zeros, in the regime where both $n$ and $N$ tend to infinity in such a way that the ratio $n/N$ converges to a positive constant $t$; we will simply write $n/N\to t$. For brevity, we will use the common abbreviations OPUC and POPUC for orthogonal polynomials on the unit circle and paraorthogonal polynomials on the unit circle, respectively, throughout the paper.
1.1. Informal summary of main results
Let us first explain the nature of our main results without introducing all necessary notation and technical points.
In our first result (Theorem 5), we establish the ratio asymptotics for polynomials $\Phi_{n,N}$, $\Phi^{*}_{n,N}$, $\Phi_{n,N}^{(\beta)}$, as $n/N\to t$, assuming the varying Nevai–López condition
$\lim_{n/N\to t}\alpha_{n,N}=\alpha.$ (1.1)
While ratio asymptotic formulas are well known in the non-varying case ($\alpha_{n,N}\equiv\alpha_{n}$), see [BarLop, Khr02] and [simon_opuc1, simon_opuc2], making the existence of such results expected, our approach via the operator theory of CMV matrices offers a streamlined proof that both extends these results to the varying case and provides a fresh insight into the classical setting. In particular, it allows for an explicit identification of the limiting ratio in terms of standard objects of OPUC theory such as the Schur, Carathéodory, and $m$-functions, and it establishes uniform convergence on optimal domains, i.e., on arbitrary compact subsets of $\mathbb{C}$} which avoid the limiting zero set of the polynomials.
Another advantage of our approach is in its broader applicability. It allows to extend the first result, without any significant additional difficulty, to much wider generality, which we demonstrate on the asymptotically periodic setting when (1.1) is replaced by the condition
@@ 1.1. Informal summary of main results
where $p\in\mathbb{N}$ is a period. This constitutes our second result (Theorem 7). In the non-varying case, it improves upon a result of Barrios and López [BarLop], see Corollary 9.
Our next results concern the asymptotics of the zero-counting measures
[…]
where $z_{j,n,N}$ and $z^{(\beta)}_{j,n,N}$ are zeros (counted with multiplicity) of $\Phi_{n,N}$ and $\Phi_{n,N}^{(\beta)}$, respectively, and $\delta_{z}$ is the Dirac unit mass measure supported at the one-point set $\{z\}$.
Clearly, the mere conditions (1.1) or (1.2) are generally insufficient to guarantee the existence of weak limits for these measures, as they constrain only a small portion of the Verblunsky coefficients. This leads us to the so-called asymptotically locally constant setting, in which we assume that
@@
where $\alpha$ is an integrable function on the interval $[0,t]$. A model situation for this set-up is obtained by sampling a piece-wise continuous function $\alpha$ on a regular grid of $[0,t]$: $\alpha_{n,N}=\alpha\left(n/N\right)$ for $n\leq tN$ (and, for instance, $\alpha_{n,N}=0$ for $n>tN$).
@@
Our third result (Theorem 10) states that, under condition (1.4), the weak limit of the zero-counting measure $\nu^{(\beta)}_{n,N}$ of POPUC exists and equals an average measure:
$\operatorname*{w-lim\ }_{n/N\to t}\nu_{n,N}^{(\beta)}=\frac{1}{t}\int_{0}^{t}\nu_{|\alpha(s)|}{\rm d}s.$ (1.5)
@@
Here the measure $\nu_{|\alpha(s)|}$ is the equilibrium measure of a circular arc determined by $|\alpha(s)|$. If we additionally assume that $|\alpha_{n-1,N}|^{1/n}\to 1$ as $n/N\to t$, then the zero-counting measure $\nu_{n,N}$ of OPUC has the same weak limit as $n/N\to t$. This is our fourth result (Theorem 18).
@@
Finally, analogous results (Theorem 14 and Remark 20) hold in the asymptotically locally periodic setting, which assumes
$\lim_{kp/N\to s}\alpha_{kp+j,N}=\alpha_{j}(s),\,\mbox{ for {a.e.} }s\in[0,t],$
@@
where $\alpha_{1},\ldots,\alpha_{p}$ are $p$ integrable functions on $[0,t]$. In this case, the measure $\nu_{|\alpha(s)|}$ from (1.5) is replaced by the equilibrium measure of a certain finite-gap subset of $\partial\mathbb{D}$}.
1.2. Relevant literature
We benefit greatly from the comprehensive treatment of the general theory of OPUC presented in Simon’s monographs [simon_opuc1, simon_opuc2].
There are numerous works that address properties of the zeros of orthogonal and paraorthogonal polynomials on the unit circle, see, e.g., [ABMV, AlfVig, BarLop, BlaSafSim, BreSee, Cas19, CasPet, DavSim, DenSim, ErdTur, golinski_maa99, Gol02, Khr03, KK, KilNen, KilSto, Siman16, Siman20, SimZero4, MFMS06a, MFMS06b, mha-saf_jat90, NevTot, Pak, SafTot, SimZero2, simon_opuc1, simon_opuc2, SimZero1, SimZero3, SimZeroRev, SimPOPUC, Sim09, SimTot, Sto, Won07]. The literature on asymptotics of orthogonal polynomials on the unit circle is so extensive that we do not attempt to survey it here. The results most relevant for our purposes, and on which we rely, are those in  [BarLop, golinski_maa99, Khr02, mha-saf_jat90, Sim09] and [simon_opuc1, Sec. 8].
@@ 1.2. Relevant literature
A parallel study formulated for orthogonal polynomials on the real line with varying Jacobi parameters has already received significant attention. Their asymptotic zero distribution was analyzed by Kuijlaars and Van Assche [kuijlaarsvanassche_jat99]. Related results had already appeared in the Szegő-type limit formulas for Hermitian variable Toeplitz matrices due to Kac, Murdock and Szegő [kms_jrma53], later revisited by Tilli [tilli_laa98]; see also [bourget-etal_jjm18, pastur_mps96]. These works also influenced the development of locally Toeplitz sequences and their applications in numerical analysis [bar-ser_20, garoni-serracapizzano_vol1, garoni-serracapizzano_vol2].
@@
The article [kuijlaarsvanassche_jat99] had a significant impact on the orthogonal polynomial community, leading to several generalizations and extensions: to the asymptotically periodic case in [vanassche_ss99, fas-ser_03], to discontinuous limiting functions in [kuijlaars-serracapizzano_jat01], to Laurent orthogonal polynomials with varying recurrence coefficients in [coussement-vanassche_jpa05], and to variable multiple orthogonal polynomials in [coussement-etal_tams08]. Our results cover the full range of [kuijlaarsvanassche_jat99, vanassche_ss99, kuijlaars-serracapizzano_jat01] in the unit-circle setting, including the analysis of both asymptotically periodic behavior and discontinuous sampling functions.
Finally, we note that the condition (1.4) appears to be very natural from the perspective of discretizing Krein systems and one-dimensional Dirac operators [denisov_imrs06, Kre55]. The problem of understanding asymptotic zero densities in the varying setting is also closely related to questions in random matrix theory concerning global eigenvalue distribution of orthogonal polynomial ensembles on the unit circle, see, e.g., [BreOfn25, DuiKoz].
1.3. Structure of the paper
@@
The paper is organized as follows. Section 2 collects notation and the necessary background from the OPUC theory and potential theory.
@@
Section 3 derives ratio asymptotics for varying POPUC and OPUC. The asymptotically constant case is treated in Section 3.1, and the asymptotically periodic case in Section 3.2.
@@
Section 4 establishes the asymptotic zero distribution for POPUC with varying Verblunsky coefficients. We present two distinct proofs of the main result: one combining ratio asymptotics with a potential-theoretic argument in Section 4.1, and the other based on the method of moments in Section 4.2. The asymptotically periodic case is discussed in Section 4.3.
@@
Section 5 is devoted to the asymptotic zero distribution for OPUC with varying Verblunsky coefficients. A transition from POPUC to OPUC is handled via balayage in Section 5.1, and the main result is deduced in Section 5.2.
@@
Finally, Section 6 presents several illustrative examples where the limiting densities are computed explicitly or expressed in terms of classical special functions, complete with numerical plots.
@@ 2. Preliminaries
2.1. Notation
Throughout the paper we use the following notation:
  • $\mathbb{N}=\{1,2,3,\dots\}$ for positive integers;
  • $\mathbb{N}_{0}=\{0,1,2,\dots\}$ for non-negative integers;
  • ${\mathbb{D}}=\{z\in{\mathbb{C}}\mid|z|<1\}$ for the open unit disk;
  • $\overline{{\mathbb{D}}}=\{z\in{\mathbb{C}}\mid|z|\leq 1\}$ for the closed unit disk;
  • ${\partial\mathbb{D}}=\{z\in{\mathbb{C}}\mid|z|=1\}$ for the unit circle;
  • $\overline{{\mathbb{C}}}={\mathbb{C}}\cup\{\infty\}$ for the Riemann sphere;
  • $\mathop{\mathrm{supp}}\nolimits\mu$ for the support of a measure $\mu$;
  • $\mathop{\mathrm{dist}}\nolimits(z,A)$ for the Euclidean distance of a point $z$ and a set $A$ in $\mathbb{C}$}.
  • ${\mathbb{D}}^{\infty}=\{\{\alpha_{n}\}_{n=0}^{\infty}\mid\alpha_{n}\in{\mathbb{D}}\mbox{ for all }n\in\mathbb{N}_{0}\}$ for the set of sequences with elements from $\mathbb{D}$};
  • $\ell^{2}(\mathbb{N}_{0})=\{\{x_{n}\}_{n=0}^{\infty}\mid\sum_{n=0}^{\infty}|x_{n}|^{2}<\infty\}$ for the Hilbert space of complex square summable sequences;
  • $\langle\cdot,\cdot\rangle$ for the inner product on $\ell^{2}(\mathbb{N}_{0})$, linear in the second argument;
  • $e_{i}$ for the $i$-th vector of the standard basis of either ${\mathbb{C}}^{n}$ or $\ell^{2}(\mathbb{N}_{0})$ which is always clear from the context, i.e., $(e_{i})_{i}=1$ and $(e_{i})_{j}=0$ for $j\neq i$.
Other notation will be introduced by its first occurrence.
@@ 2.1. Notation
$\{n_{j}\}_{j\in\mathbb{N}},\{N_{j}\}_{j\in\mathbb{N}}\subseteq\mathbb{N},\mbox{ such that }n_{j},N_{j}\to\infty\mbox{ and }n_{j}/N_{j}\to t,\mbox{ as }j\to\infty.$ (2.1)
@@
expresses the limit of the doubly indexed sequence of measures $\mu_{n,N}$ converging to $\mu$ weakly, i.e., in the weak topology.
@@ 2.2. Orthogonal and paraorthogonal polynomials
Given a non-trivial probability measure $\mu$ on $\partial\mathbb{D}$}, i.e., one with infinite support, let $\Phi_{n}$ be the unique monic polynomial of degree $n$ satisfying the orthogonality relation
$\int_{0}^{2\pi}\Phi_{n}({\rm e}^{{\rm i}\theta}){\rm e}^{-{\rm i}k\theta}\,{\rm d}\mu({\rm e}^{{\rm i}\theta})=0,\quad\mbox{for all }k=0,\ldots,n-1.$
These OPUC satisfy the Szegő recursion relations:
@@
If the Verblunsky coefficient $\alpha_{n-1}$ is replaced by a unimodular parameter $\beta\in{\partial\mathbb{D}}$, one obtains the POPUC: